A nurse is to administer neomycin sulfate (Mycifradin) 1g po BID. Available is a suspension containing 500 mg per teaspoon. How many tablespoons should a nurse administer each week
Answer #1

Ordered dose of Mycifradin= 1 g

Available dose on hand = 500 mg / tsp

i) conversion:

1 gram = 1000 mg

ii) Dose for patient:

If 1 teaspoon contains 500 mg  

Then X teaspoon contains 1000 mg

By cross multiplication,

therefore ;

1 \times 1000 = 500 \times X

Therefore X = 1000 \div 500 = 2

Therefore patient will get 2 teaspoon in single dose dose.

iii ) But patient is ordered BID dose.

Therefore patient will get total 4 teaspoon in a whole day.

There are total 7 days in a week.

Therefore patient will get total 7 \times 4 =28 teaspoon in a week.